40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11099     Entry time: Thu Mar 5 04:29:13 2015
Author: ericq 
Type: Update 
Category: LSC 
Subject: Locking work tonight 

Brief elog of my activities tonight:

I was able to transition the digitial CARM control to REFL11 through the common mode board a total of one time, lock broke after a few seconds.

My suspicion was that when we did this on Monday, we unintentionally had a reasonable DARM offset, which reduced the finesse enough to let us take linear transfer functions and hop over. So, tonight, I intentionally looked at transitioning to CM_SLOW at some DARM offset. Using DARM offset of a few times 0.1 really calms the "buzzing" down, and makes it fairly straightforward to measure linear CARM sensing TFs. However, the CARM optical plant seems to change a fair amount depending on the DARM offset, in such a way that I was not able to compensate well enough to repeatedly transition.


Before I did anything else tonight, I measured the ALS noise down to 0.1 Hz, as a benchmark of how things are behaving.

With the arms locked on POX/POY, the HZ calibrated ALS channels reported

  • ALSX : 471Hz RMS
  • ALSY: 298 Hz RMS

Then, with the arms CARM/DARM locked on ALS, the PDH signals reported (using a line and the HZ channels for conversion)

  • Xarm : 552 Hz RMS
  • Yarm : 264 Hz RMS

Not bad! I roughly estimate this to mean ~90pm RMS CARM/DARM motion. (If X was as good as Y, it would be ~50pm...)


Some things I feel are worth noting:

  • In an effort to avoid the ETMX issues that Jenne had last night. I used MC2 to actuate CARM, and 2xETMY to actuate DARM. None of my locklosses appeared to be due to saturation of DARM, so I think it worked fine. The main drawback seems to be that if you have a violent lock loss, you may have to wait a bit for the IMC to relock; this only happened once tonight.
  • After the IR resonance finding scripts, I would run a z servo to try and get the PDH signal to cross zero. This made the ALS CARM and DARM zeros closer to the real resonating zeros than I usually see.
  • It is lately possible to sit at higher powers (albeit with very high RIN) for sizable amounts of time. In my last lock, I was in the range of 10-60x single arm power for around 30 minutes before I blew it with a failed transition attempt.
  • The set points for the QPD servos don't change much from lock to lock. I didn't have any problem using them tonight.

Tomorrow, I'll post some transfer functions of the difference between the ALS and CARM plants that I measured.

ELOG V3.1.3-